A tank of volume 0.02 m^3 contains 0.20 kg of helium at 27 oC.

(a) How many moles of helium are present? (hint: the relative atomic mass in grams is one mole)

(b) Calculate the pressure of the gas in Pa. The universal molar gas constant R = 8.31 J K–1 mol–1


1
Expert's answer
2021-03-31T09:39:39-0400

(a) "\nu=\frac{m}{M}=\frac{0.2}{4\cdot10^{-3}}=50 \ (mole)"


(b) "pV=\nu RT\to p=\frac{\nu RT}{V}=\frac{50\cdot8.31\cdot(273+27)}{0.02}=6232500\ (Pa)"